OpenStudy (anonymous):

how to solve (x+2)^3/4=27

OpenStudy (anonymous):

-(x+2)^-3/4=-1/27 ......1 - --------------= -1/27 ..∜(x+2)^3 ................1 1/27 = ----------- ..........∜(x+2)^3 (∜(x+2)^3)^4 = 27^4 (x+2)^3 = 531,441 x+2 = ∛(531,441) x= -2 +81 x= 79 answer//

OpenStudy (anonymous):

\[\large{(x+2)}^{\frac{3}{4}}= 27\] \[\large{(x+2)} = 27^{\frac{4}{3}}\] \[x +2 = 3^4\] \[x = 81-2\] \[x = 79\]

OpenStudy (anonymous):

I thought it was\[\frac{(x+2)^3}{4} = 27\]

hero (hero):

I like Ishaan's solution better

OpenStudy (anonymous):

\[x+2=\pm27^{\frac{4}{3}}\] \[x+2=\pm\sqrt[3]{27^4}\] \[x+2=\pm3^4\] \[x+2=81\] \[x=79\] or \[x+2=-81\] \[x=-83\]

OpenStudy (anonymous):

i think in these cases you get two solutions

OpenStudy (anonymous):

but it is 1/3 power I don't think they have plus and minus like square root right ? I am not sure

OpenStudy (anonymous):

on no i am wrong. it is \[\sqrt[4]{(x+2)^3}\] sorry

OpenStudy (anonymous):

can't take the fourth root of negative. ignore me.